# spacing=nonfrench

### v1

@@ Appendix A Advanced Sequence Number Manipulation
We verified all techniques successfully against PuTTY 0.79. Additionally, our experiments show that OpenSSH 9.5p1 recognizes a rollover of sequence numbers and terminates the connection, thus not being affected by any advanced technique. AsyncSSH 2.13.2 and libssh 0.10.5 terminate the connection due to handshake timeouts before these techniques conclude. Dropbear 2022.83 disconnects on Unknown messages instead of responding with Unimplemented, but allows $\mathsf{Rcv}$ to roll over, therefore being affected by RcvDecrease only.
(a) RcvIncrease
(b) RcvDecrease
(c) SndIncrease
(d) SndDecrease
Appendix B More Extensions
These extensions are defined in addition to those listed in Section 5.
  • no-flow-control - Negotiable extension to disable flow control in the SSH connection layer [9].
  • delay-compression [9] - Negotiable extension to renegotiate compression algorithms as delayed compression without a key re-exchange. Delayed compression in the context of SSH refers to any compression that will take effect after successful client authentication.
  • elevation - Informational client-side extension to specify the desired security context in case of administrator logins in Windows operating systems [9].
  • global-requests-ok [10] - Informational extension to announce global requests within the SSH connection layer will be handled in accordance with RFC 4254.
  • ext-auth-info [8] - Client-side extension to signal support for extended authentication failure messages during client authentication containing an arbitrary set of key-value pairs.

### v2

@@ A.4.1 Major Claims
Sequence Number Manipulation (Sect. 4.1). We verified all techniques successfully against PuTTY 0.79. Additionally, our experiments show that OpenSSH 9.5p1 recognizes a rollover of sequence numbers and terminates the connection, thus not affected by any technique but RcvIncrease. AsyncSSH 2.13.2 and libssh 0.10.5 allow for RcvIncrease but terminate the connection due to handshake timeouts before any advanced technique concludes. Dropbear 2022.83 disconnects on Unknown messages instead of responding with Unimplemented but allows $\mathsf{Rcv}$ to roll over, therefore being affected by RcvIncrease and RcvDecrease only.
(C2):
Extension Downgrade (Sect. 5.2). We successfully evaluated the attack in 10,000 trials on ChaCha20-Poly1305 and CBC-EtM against OpenSSH 9.5p1 and PuTTY 0.79 clients, connecting to OpenSSH 9.4p1 (Unknown only) and 9.5p1. For CBC-EtM, our success rate in practice was 0.0003 (OpenSSH) resp. 0.0300 (PuTTY), improved to 0.0074 (OpenSSH) resp. 0.8383 (PuTTY) when sending Ping instead of Unknown.
(C3):
Rogue Extension Negotiation (Sect. 6.1). We successfully evaluated the attack against AsyncSSH 2.13.2 as a client, connecting to AsyncSSH 2.13.2.
(C4):
Rogue Session Attack (Sect. 6.2). We successfully evaluated the attack against AsyncSSH 2.13.2 as a server, connecting to AsyncSSH 2.13.2.
Internet Scan (Sect. 7)
We are also including sample data, aggregated data, and evaluation scripts on the Internet scan.
A.4.2 Experiments
The evaluation scripts (in the directory scripts) are interactive and self-describing. Some of them have several output files. In that case, the files (as described below) are all opened in the text file viewer less at the same time, requiring keyboard-based navigation to see all of the results. As a gentle introduction to less, see Table 6 for a quick reference of useful keyboard shortcuts.
Shortcut Description
q Quit
h Help
S Wrap long lines on/off
/ Search
:n Next file
:p Previous file
Table 6: Common keyboard shortcuts of less.
(E1):
test-sqn-manipulation.sh [$\approx 1-3$ hours per client/variant combination]: Run one of the four sequence number manipulation attacks to prove (C1). RcvIncrease is very fast; the others can be slow.
Execution:
After starting the script, choose a client, one of the four attack options, and input the manipulation offset $N$. To prove (C1), input $N=1$.
Results:
The attack is complete once the progress bar fills. After that, there will be an error message because the secure channel is broken, as the script does not implement any prefix truncation to complete the attack.
(E2a):
test-ext-downgrade.sh [$\approx 1$ minute]: Run the extension downgrade attack to prove (C2) for ChaCha20-Poly1305.
Execution:
After starting the script, choose an arbitrary client and server combination. Afterward, choose attack variant 1 to select ChaCha20-Poly1305.
Results:
The script will conclude by opening the following files simultaneously in less:
  1. 1.
    diff of files 3 and 4
  2. 2.
    diff of files 5 and 6
  3. 3.
    Server log (unmodified connection)
  4. 4.
    Server log (tampered connection)
  5. 5.
    Client log (unmodified connection)
  6. 6.
    Client log (tampered connection)
  7. 7.
    PoC proxy log
Navigate to the second file. The file compares the output of the selected SSH client in the case of an extension downgrade attack to the output of an unmodified connection. The diff will indicate the presence of SSH_MSG_EXT_INFO and absence of SSH_MSG_IGNORE in the unmodified connection only, thus proving (C2) for ChaCha20-Poly1305.
(E2b):
bench-ext-downgrade.sh [$\approx 1-2$ hours per client/variant combination]: Run the extension downgrade attack 10,000 times to prove (C2) for CBC-EtM (Unknown and Ping).
Execution:
After starting the script, choose between Unknown and Ping variants of the attack, then select between OpenSSH and PuTTY client. A progress bar will show the current trial.
Results:
After finishing all trial connections, the number of successful trial runs will be outputted to the console. The relative success rate will be close to the values claimed in (C2), thus proving the functionality and success probability claims in (C2) in the case of CBC-EtM.
(E3):
test-asyncssh-rogue-ext-negotiation.sh [$\approx 1$ minute]: Run rogue extension attack to prove (C3).
Execution:
The attack is automatic.
Results:
The script will conclude by opening a set of seven files in less. Refer to the results of (E2a) for a list of files opened. Navigate to the second file. The diff will indicate the presence of the server-sig-algs extension with an attacker-chosen value in the tampered connection, thus proving (C3).
(E4):
test-asyncssh-rogue-session-attack.sh [$\approx 1$ minute]: Run the rogue session attack to prove (C4).
Execution:
The attack is automatic.
Results:
The script will conclude by opening a set of seven files in less. Refer to the results of (E2a) for a list of files opened. Navigate to the first file. The diff will indicate successful authentication for the victim (unmodified connection) and attacker (tampered connection), respectively. Afterward, navigate to the second file and examine the output of each client connection at the end of the file. In the unmodified connection, the server will respond with the username victim, while in the attacked connection, the server will respond with the username attacker. This proves (C4).
(E5):
scan_util.py [$\approx 1$ minute]: Run the script to aggregate a set of zgrab2 scan results. Note that this script is in the sub-directory scans. Without the full data, we can not prove the statistics in our paper. However, we can demonstrate how we aggregated the scan results and classified the algorithms.
Execution:
Build the docker image by running the following command inside the scans directory:
1\$ docker build . -t terrapin-artifacts/scan-util
Now aggregate the sample.json file which can be found in the sample sub-directory by running the following command inside the scans directory:
1\$ docker run --rm -v ./sample:/input terrapin-artifacts/scan-util evaluate -i /input/sample.json -o /input/sample-ae.acc.json
Results:
The aggregation result will become available as sample-ae.acc.json inside the sample sub-directory. The total number of clients, status and version distribution, offered key exchange algorithms, comment strings, and other evaluation criteria will match the data present within the sample.json file. Also, there is no difference between the sample-ae.acc.json and sample.acc.json files (aside from the evaluation start and end timestamps).
A.4.3 Troubleshooting
Address already in use.
If an attack script is interrupted, some docker containers may not be cleaned up properly, blocking the server port permanently or for the duration of TIME-WAIT (1 min. on Linux, 30 sec. on MacOS).
Please follow these steps in this case:
  1. 1.
    Run the script cleanup-system.sh. This will stop and remove any pending Docker containers.
  2. 2.
    If the problem persists, wait for up to 4 minutes.
System Reset.
To fully clean up the Docker containers and images, you can run cleanup-system.sh --full.
A.5 Notes on Reusability
The proof-of-concept code (pocs/) has been kept short for simplicity and is thus not modularized for reusability. However, the artifacts may serve as a template for other MitM attacks on network protocols like SSH.
The Docker files for the evaluated SSH implementations (impl/) may be generally useful in other research on SSH.
Improvements to Wireshark for better dissection of SSH protocols (not included in these artifacts) have been submitted and accepted upstream and will be available in a future version of Wireshark.
A.6 Version
Based on the LaTeX template for Artifact Evaluation V20231005. Submission, reviewing and badging methodology followed for the evaluation of this artifact can be found at https://secartifacts.github.io/usenixsec2024/.